OpenStudy (anonymous):

given the following quadratic inequality solve. x^2+17x+72>0

OpenStudy (ali2x2):

Raise x to the 2nd power

OpenStudy (ali2x2):

(x2 + 17x) + 72 > 0

OpenStudy (ali2x2):

Simplify x2+17x + 72

OpenStudy (ali2x2):

Factor x2+17x+72

OpenStudy (ali2x2):

The first term is, x2 its coefficient is 1 . The middle term is, +17x its coefficient is 17 . The last term, "the constant", is +72

OpenStudy (anonymous):

is it x<-9 and x>-8 or is it x<-9 or x>-8?

OpenStudy (ali2x2):

Step-1 : Multiply the coefficient of the first term by the constant 1 • 72 = 72 Step-2 : Find two factors of 72 whose sum equals the coefficient of the middle term, which is 17 .

OpenStudy (ali2x2):

-72 + -1 = -73 -36 + -2 = -38 -24 + -3 = -27 -18 + -4 = -22 -12 + -6 = -18 -9 + -8 = -17 -8 + -9 = -17 -6 + -12 = -18 -4 + -18 = -22 -3 + -24 = -27 -2 + -36 = -38 -1 + -72 = -73 1 + 72 = 73 2 + 36 = 38 3 + 24 = 27 4 + 18 = 22 6 + 12 = 18 8 + 9 = 17 That's it

OpenStudy (ali2x2):

Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 8 and 9 x2 + 8x + 9x + 72

OpenStudy (ali2x2):

Step-4 : Add up the first 2 terms, pulling out like factors : x • (x+8) Add up the last 2 terms, pulling out common factors : 9 • (x+8)

OpenStudy (ali2x2):

Step-5 : Add up the four terms of step 4 : (x+9) • (x+8) Which is the desired factorization

OpenStudy (ali2x2):

(x + 9) • (x + 8) > 0

OpenStudy (ali2x2):

Solve (x+9)•(x+8) > 0

OpenStudy (ali2x2):

there

OpenStudy (anonymous):

i know i solved it i just need to know if its x<-9 AND x>-8 or is it x<-9 OR x>-8?

OpenStudy (ali2x2):

oh

OpenStudy (ali2x2):

well my answer tells u c;brb